Controls Meeting 21-Mar-2013

From GlueXWiki
Jump to: navigation, search


  • Announcements
  • Minutes from previous meeting 7-Mar-2013
  • Requirements for BCAL/FCAL LED pulser calibration system - Elton
  • Update on FCAL dark room - Dan, John, Manuel
  • Update on PLC test stand - Dave
  • Counting House status - Hovanes, Elliott
    • furniture, computers, networks, etc.
  • EPICS update - Hovanes
  • Solenoid update - Mark, Scot, Elliott
  • Progress on other controls applications - Dave
    • storing/retrieving PLC parameters
  • What is the role of the JInventory system - All
    • relation to DAQ, HV/LV control, cabling?
    • need to finalize conventions and get all equipment entered
    • need to purchase/print and install barcode labels
    • some software improvements and bug fixes remain
    • other halls interested, may provide programming manpower
  • Mantis task tracker
    • start using this again?
  • Midas elog

  • Experiment Control Supervisor - Elliott


9:30AM Thurs 21-Mar-2013 CC A110


Next Meeting

9:30am Thurs 4-Apr-2013


Present: Elliott W, Elton S, Dan B, Fernando B, John L, Yi Q, Scot S, Mark S, Dave B, Tara B, Alex S, Hovanes E.


Elton gave a thorough overview of the LED pulser calibration system:

  • System needs to send trigger to TS and TTL pulses to LED flashers.
  • External power includes 5V and bias to set LED intensity, will use SiPM power supplies.
  • LED flashed via capacitor discharge.
  • Will have LED boards, distribution boards and external TTL pulsing system.
  • LED boards done, distribution boards and pulsing system to be designed.
  • For trigger timing all cables will be equal length, will get ~5ns jitter.
  • Need 50' LEMO cables from VME crates to distribution cards.
  • BCAL needs 32 TTL channels going to 8 distribution boards, 4 per end.
  • Cables from distribution boards about 12'.
  • 4 lines go to each LED board.
  • Will use three CAEN V1495 modules to generate TTL pulses, about $5k each incl. mezzanine boards.
  • Max 16 TTL channels per V1495.
  • Will likely use two VME crates, one for BCAL with two V1495's, one for FCAL with one V1495.
  • Not clear if crates need to be purchased, hope to find free slots in other crates.
  • Hope to not have to put V1495's in DAQ crates (discriminator crates are ok).
  • V1495 has many LVDS outputs suitable for input to TS.
  • FCAL has 10 LED boards per quadrant, 4 LED's per board, three colors.
  • FCAL does not use distribution board, TTL signals go directly to LED boards.
  • Ben will program V1495, 2 man-weeks allocated.
  • Need to program pulse width, trig-to-pulse delay, start, stop, #pulses, etc.
  • Once written and tested the FPGA code is not expected to need to change.
  • V1495 has programmable switches that could be used to implement local operational modes.
  • Will use EPICS to control V1495 operation.
  • EPICS driver needs to be written (Nerses or Hovanes).
  • Considering running pulsers all the time at low rate during data taking.
  • Max conceivable pulse rate 1 MHz, less ok.

FCAL Controls

John and Dan reported:

  • Dark room parts list to be sent out, want feedback from JLab techs on choices.
  • If no comments received will choose parts themselves.
  • Dark room might be done in 2-3 months.
  • One working power supply here, remainder in production or repair.
  • Will soon have enough equipment to test FCAL in manual mode.

Counting House

  • Rack room is ready and working, some computer and network equipment will be ordered in May.
  • Furniture ordered, should arrive in a few weeks.
  • Some conference equipment not ordered yet.
  • Chairs finally ordered.

PLC Test Stand

  • Parts coming.
  • Need to test EPICS connection to CompactLogix processor (SNS swears it works).
  • Dave will create tags, Hovanes will try to access them via EPICS IOC.


  • Beni and Hovanes working on auto-configuring power supply drivers.
  • Nerses may finally be coming in May!


Still waiting for cryo plant to continue cooldown. Yet again they had to warm up to clean out carbon contamination. Power supply still being repaired, Accelerator helping to diagnose problem. Succeeded in ramping to 1500A with just one pass bank, transistors in second bank being replaced. Dave L is working out mapping granularity. George borrowed two Hall probes.

PLC Controls

Still collecting requirements from detector groups. Target understood, testing gas system components. Dave will discuss with Jonathan whether we need to implement a backup/restore system for the solenoid test (using BURT, an EPICS utility). Eventually Accelerator will implement restore functionality in MyA, their archiver. Dave will think about how to group tags into useful subsets for backup/restore purposes.


Elliott noted that the time has come to seriously think about the role of the inventory system in the experiment w/r/to translation tables, HV/LV control, etc. Also need to complete entry of equipment and to barcode everything.


Elliott wants to begin using this again. He will bring it up to date.